# WarriorCoder: Learning from Expert Battles to Augment Code Large Language Models

### v1

@@ Abstract
Despite recent progress achieved by code large language models (LLMs), their remarkable abilities are largely dependent on fine-tuning on the high-quality data, posing challenges for data collection and annotation. To address this, current methods often design various data flywheels to gather complex code instructions, enabling models to handle more intricate tasks. However, these approaches typically rely on off-the-shelf datasets and data augmentation from the limited pool of proprietary LLMs (e.g., Claude, GPT4, and so on), which limits the diversity of the constructed data and makes it prone to systemic biases. In this paper, we propose WarriorCoder which learns from expert battles to address these limitations. Specifically, we create an arena for current expert code LLMs, where each model challenges and responds to others’ challenges, with evaluations conducted by uninvolved judge models. This competitive framework generates novel training data constructed from scratch, harnessing the strengths of all participants. Experimental results demonstrate that WarriorCoder achieves competitive performance compared to previous methods, even without relying on proprietary LLMs.
@@ 1 Introduction
Figure 1: The comparisons between our method and traditional data flywheels. Different from the previous work, we create an arena for current expert code LLMs. No demand for seed datasets, human prompts and proprietary models’ annotations, the target model learns from the pairwise competitions and integrates the strengths of competitors.
@@
To address these challenges, some approaches propose various data flywheels to generate instruction data. Building on Self-Instruct, Chaudhary (2023) constructs Code Alpaca by prompting teacher LLMs to generate instructions in a few-shot setting. To further enhance the diversity and complexity of Code Alpaca, WizardCoder Luo et al. (2024b) employs Evol-Instruct to evolve the original instructions. These methods apply general data augmentation to instruction construction, lacking specific design considerations for the code domain. Given that, recent methods specifically design frameworks for instruction generation tailored to code. For example, WaveCoder Yu et al. (2024) collects raw code snippets and defines different tasks based on them. Similarly, InverseCoder Wu et al. (2024b) prompts code LLMs to generate high-quality instructions based on the original code through techniques like code summarization and self-evaluation. These methods construct data for code in various ways, effectively enhancing the model’s code generation capabilities. However, they still rely on existing datasets Muennighoff et al. (2024) and calls to proprietary LLMs (e.g., GPT-3.5, GPT-4, etc.), making data collection costly. Additionally, the limited data sources and annotators constrain the diversity Yu et al. (2023); Wang et al. (2023) of the data and inherit the system biases inherent in the limited pool of annotators Wei et al. (2024).
@@
The challenges mentioned above have motivated us to propose WarriorCoder which learns from expert battles to address the current limitations. Specifically, we create an arena for current expert code LLMs and organize competitions among them. As shown in Figure 1, the models compete against each other, with evaluations conducted by uninvolved models. The attacker challenges all other models in its area of expertise, and the target model learns from the winner of these pairwise competitions. In this way, we can generate novel training data that incorporates the strengths of all the expert LLMs. Different from the previous work, our training data is fully automatically constructed from scratch, rather than expanding available datasets by prompting limited proprietary LLMs. In essence, the target model absorbs the best practices from a variety of powerful code LLMs, enabling it to be theoretically extended indefinitely to include emerging expert code LLMs. Moreover, our method eliminates the need for human involvement and proprietary LLMs in the training process, making it possible to collect high-quality, diverse data at a low cost. The main contributions of this paper are summarized as follows:
  • We identify the limitations of current data flywheels through detailed analysis and propose WarriorCoder, a new paradigm in which the target model absorbs best practices from a variety of strong code LLMs, rather than learning from existing datasets or their expansions.
  • We create an arena for current strong LLMs and generate novel training data that integrate the strengths of all participants, which can be easily scaled to accommodate the growth of expert code LLMs.
  • Extensive experiments demonstrate the effectiveness of WarriorCoder, with ablation and comparison studies explaining how and why it works. It achieves competitive performance compared to previous methods, even without relying on proprietary LLMs.
@@ 2.1 Code LLMs
Code plays a crucial role in application areas for LLMs Jain et al. (2024), attracting significant interest from both academia and industry. Codex Chen et al. (2021), an LLM with 12 billion parameters, can solve 72.31% of complex Python programming problems posed by humans. Following the success of Codex Lyu et al. (2024), the rise of new code LLMs has demonstrated even greater capabilities, such as code generation and debugging, as model sizes continue to grow Hou et al. (2023). These models have become indispensable in a variety of code-specific tasks Zan et al. (2023). Despite this impressive progress, the performance of current open-source models still lags behind that of proprietary ones (e.g., GPT-3.5, GPT-4, etc.), primarily because stronger models often keep their training data proprietary Hui et al. (2024). As a result, the lack of publicly available code datasets remains a significant barrier to further development in this field.
@@ 3 WarriorCoder: Learning from Expert Battles
Figure 2: The diagram of learning from a battle between two chatbots.
In this section, we describe how WarriorCoder learns from expert battles. Different from the previous work, we build an arena for current expert code LLMs and organize competitions among them, other than expanding existing datasets through prompting the limited pool of proprietary LLMs. Each model takes advantage of what he has learned to challenge others under the evaluation of uninvolved judge models. The target model learns from the winner of these pairwise competitions and integrates the strengths of all the competitors.
@@ 3.1 Instruction Mining from Scratch
Considering a battle between A and B where A is on offense and B is on defense. The first step of our method is to use A’s strengths to challenge B, which makes it necessary to know what A has learned during its training process. However, nearly all the current open-sourced LLMs keep their core data proprietary. Inspired by Magpie Xu et al. (2024b), we design a completion-based method to mine the capabilities which the code LLMs have already mastered. Here we take Qwen2.5 Hui et al. (2024) as an example. A conversation about writing Python code in the chat template of Qwen2.5 is:
@@
Such chat templates are pre-defined conversational structures to guide the interaction between the model and the user. Based on the LLMs’ strong completion abilities, we feed only the prefix of the chat template into them, prompting the LLM to generate the user queries:
@@
In this way, we can collect various instructions the model has already learned under various generation settings (different values of temperature and top-p). Unlike traditional synthetic data, our instructions are not synthesized by the models but directly sampled from their distributions, which avoids pattern overfitting and significant shifts in the output distribution Chen et al. (2024b). However, these instructions may be repetitive, ambiguous, unclear, or too easy. To address these concerns, we first deduplicate the data and adopt uninvolved LLMs in the arena to assess their difficulty. We divide the difficulty of instructions into four levels:
  • Excellent (9-10): For queries that are very clear, specific, and well-articulated. These queries are particularly challenging and excellently designed to assess the AI’s proficiency.
  • Good (6-8): For queries that are clear and specific instructions. These are not overly difficult to answer and moderately assess the AI’s capabilities.
  • Average (3-5): For queries that are fairly clear and specific instructions. These queries are easy to answer.
  • Poor (1-2): For queries that are ambiguous or unclear.
Only good and excellent instructions are considered during the following steps (the top half of Figure 2).
Then we compress the high-quality instructions for the efficiency of post-training. To ensure the diversity and representativeness of instructions, we employ KCenterGreedy algorithm Sener and Savarese (2018) to select the final instructions based on the embedding model - all-roberta-large-v1 Liu et al. (2019).
@@ 3.2 Win-Loss Decision
The defender is required to respond to the attacker’s question, while the attacker A must also provide an answer to its own query. Once both responses are collected, we use other LLMs in the arena to evaluate the correctness and helpfulness of the pairwise data (the bottom half of Figure 2). To ensure impartiality in the evaluation process, we establish a set of rules, including order shuffling, suspicion averting, and offensive-defense balance.
Having the voting results, we can calculate the local score for each query $q$:
$x^{q}_{A>B}=\frac{t_{A}}{t_{A}+t_{B}}\quad x^{q}_{B>A}=\frac{t_{B}}{t_{A}+t_{B}}$ (1)
where $x^{q}_{A>B}$ and $x^{q}_{B>A}$ are A’s and B’s local score for $q$. $x^{q}_{A>B}$ represents the percentage of votes that candidate A receives, while $x^{q}_{B>A}$ similarly represents the percentage of votes that candidate B receives. $t_{A}$ and $t_{B}$ are the number of votes which A and B win.
@@
where $X^{Elo}_{A>B}$ and $X^{Elo}_{B>A}$ indicate the expected probabilities of A defeating B and B defeating A, respectively. $R_{A}$ and $R_{B}$ are the Elo rating of A and B, which are updated dynamically and iteratively. Given the battle result of A and B on an instruction $q$, we update them by:
$R_{A}\leftarrow R_{A}+K\times(s^{q}_{A>B}-X^{Elo}_{A>B})$ (3)
$R_{B}\leftarrow R_{B}+K\times(s^{q}_{B>A}-X^{Elo}_{B>A})$
where $s^{q}_{A>B}$ and $s^{q}_{B>A}$ are the actual scores of player A and B (1 for a win, 0.5 for a draw, and 0 for a loss). The factor $K$ controls the sensitivity of rating changes. Here we set it to 0.005.
Based on Equation 1 and Equation 2, we can obtain the final score of A’s response for instruction $q$:
$e^{q}_{A}=\sum_{B\in Com\setminus A}\alpha X^{Elo}_{A>B}+(1-\alpha)x^{q}_{A>B}$ (4)
where $Com$ is the set of all the competitors and $\setminus$ is the subtraction operation. $\alpha$ is the coefficient to balance the local contingency and global consistency.
@@ 3.3 Final Training
Each item of the constructed dataset consists of an instruction, responses from various strong LLMs, and their corresponding scores, which supports multiple post-training methods like SFT and KTO Ethayarajh et al. (2024). It’s worth noting that we employ KTO instead of DPO Rafailov et al. (2023) because KTO utilizes all the responses while DPO only makes use of pairwised data. We select the response with the highest score as the gold output for SFT and set a threhold to distinguish the chosen and rejected completions for KTO.
@@ 4 Experiments
4.1 Experimental Details
Backbones
We use DeepSeekCoder-Base-6.7B Guo et al. (2024) to initialize the target model. As for the competitors of expert battles, we choose strong open-source LLMs including Llama-3.1 Dubey et al. (2024), Qwen2.5 Hui et al. (2024), and Mistral Jiang et al. (2023).
Datasets
To evaluate the effectiveness of our model, we conduct evaluations on HumanEval Chen et al. (2021), MBPP Austin et al. (2021), HumanEval+ Liu et al. (2023), and MBPP+ Liu et al. (2024). For a fair comparison, we use greedy decoding for generating answers to each question, consistent with previous work.
Experimental Settings
We conducted parallel training on 8 NVIDIA A800 80G GPUs for all backbone models. The global batch size was set to 512, and the number of epochs was set to 2. We used a learning rate of $1\times 10^{-5}$ and a weight decay of $3\times 10^{-7}$. Additionally, a WarmupLR scheduler with a warmup ratio of 0.2 was used. For both the offensive and defensive turns, the number was set to 70,000.
HumanEval HumanEval+ MBPP MBPP+ Rely on proprietary LLMs?
@@ 4.1 Experimental Details
Ours WarriorCoder-DS (6.7B) 72.1 (+24.5) 66.0 (+26.4) 70.5 (+0.3) 57.8 (+1.2)
Table 1: The pass@1(%)on four benchmarks.
@@ 4.2 Main Results
The overall results are summarized in Table 1. WarriorCoder achieves impressive performance, with a pass@1 rate of 72.1% (66.0%) on HumanEval (HumanEval+) and 70.5% (57.8%) on MBPP (MBPP+), surpassing nearly all other fine-tuned models. Notably, it shows a significant boost over the backbone model, particularly on HumanEval and HumanEval+ (with improvements of 24.5 and 26.4, respectively). WarriorCoder also outperforms both Magicoder-DS and WaveCoder-DS, which share the same backbone architecture and similar amounts of training data. This substantial performance gap highlights the effectiveness of our approach in generating higher-quality training data, providing a clear advantage over other models that rely on similar foundational setups.
Moreover, the performance of WaveCoder-DS on MBPP drops to 63.6% after fine-tuning, while our model improves to 70.5%. These results underscore WarriorCoder’s paradigm as a powerful data flywheel that effectively absorbs abilities from various expert code LLMs. Our method significantly enhances the capabilities of target models to generalize across the four benchmarks, showcasing its superiority in utilizing diverse data sources to drive performance improvements.
@@
Moreover, previous data flywheels typically rely on augmentations and annotations generated using proprietary large language models (LLMs) and specially designed prompts. In contrast, our approach does not require human-generated prompts or proprietary LLMs. Despite this, we achieve competitive results that rival those of advanced code LLMs that depend on such resources. This underscores the quality of our data flywheel, demonstrating its capacity to significantly enhance the target model at a much lower cost.

### v3

@@ Abstract
Despite recent progress achieved by code large language models (LLMs), their remarkable abilities are largely dependent on fine-tuning on the high-quality data, posing challenges for data collection and annotation. To address this, current methods often design various data flywheels to collect complex code instructions, enabling models to handle more intricate tasks. However, these approaches typically rely on off-the-shelf datasets and data augmentation from a limited set of proprietary LLMs (e.g., Claude, GPT4, and so on), which restricts the diversity of the constructed data and makes it prone to systemic biases. In this paper, we propose WarriorCoder, a novel paradigm learns from expert battles to address these limitations. Specifically, we create an arena where leading expert code LLMs challenge each other, with evaluations conducted by impartial judges. This competitive framework generates novel training data from scratch, leveraging the strengths of all participants. Experimental results show that WarriorCoder achieves state-of-the-art performance compared to previous models of the same size, even without relying on proprietary LLMs.
@@ 1 Introduction
Figure 1: The comparisons between our method and traditional data flywheels. Unlike previous work, we guides the target model to learn from pairwise competitions. No demand for seed datasets, human-generated prompts, or annotations from proprietary models, the target model integrates the strengths of its competitors.
@@
To address these challenges, some approaches propose various data flywheels to generate instruction data. Building on Self-Instruct, Chaudhary (2023) constructs Code Alpaca by prompting teacher LLMs to generate instructions in a few-shot setting. To further enhance the diversity and complexity of Code Alpaca, WizardCoder Luo et al. (2024b) employs Evol-Instruct to evolve the original instructions. These methods apply general data augmentation to instruction construction, lacking specific design considerations for the code domain. Given that, recent methods specifically design frameworks for instruction generation tailored to code. For example, WaveCoder Yu et al. (2024) collects raw code snippets and defines different tasks based on them. Similarly, InverseCoder Wu et al. (2024b) prompts code LLMs to generate high-quality instructions based on the original code through techniques like code summarization and self-evaluation. These methods construct data for code in various ways, effectively enhancing the model’s code generation capabilities. However, they still rely on existing datasets Muennighoff et al. (2024) and calls for proprietary LLMs (e.g., GPT-3.5, GPT-4, etc.), making data collection costly. Additionally, the limited data sources and annotators constrain the diversity Yu et al. (2023); Wang et al. (2023) of the data and inherit the system biases inherent in the limited pool of annotators Wei et al. (2024).
@@
The challenges mentioned above motivated us to propose WarriorCoder, which learns from expert battles to overcome current limitations. As illustrated in Figure 1, the attacker challenges the opponent within its area of expertise, and the target model learns from the winner of these pairwise competitions. Specifically, we design a completion-based method to mine the capabilities which the attacker has already mastered, then integrate Elo Rating and voting results to balance the local and global evaluation. This approach enables us to generate novel training data from scratch, incorporating the strengths of all the expert code LLMs, rather than relying on limited proprietary LLMs to expand existing datasets. Moreover, our method eliminates the need for human involvement and proprietary LLMs in the data collection, making it possible to collect high-quality, diverse data at a low cost. The main contributions of this paper are summarized as follows:
  • We identify the limitations of current data flywheel and propose a new scalable paradigm where the target model learns from expert battles to solve them.
  • We design a completion-based method for collecting instructions and introduce the Elo Rating system for evaluating responses, enabling the creation of high-quality and diverse training data at a low cost. Fine-tuned on this data, WarriorCoder incorporates the strengths of all the experts, achieving state-of-the-art performance compared to previous models of the same size, without relying on proprietary LLMs.
  • Extensive experiments demonstrate the excellent performance of WarriorCoder on multiple code-related tasks, with ablation and analysis studies explaining how and why it works.
@@ 2.1 Code LLMs
Code plays a crucial role in application areas for LLMs Jain et al. (2024), attracting significant interest from both academia and industry. Codex Chen et al. (2021), an LLM with 12 billion parameters, can solve 72.31% of complex Python programming problems. Following the success of Codex Lyu et al. (2024), the rise of new code LLMs has demonstrated even greater capabilities, such as code generation and debugging, as model sizes continue to grow Hou et al. (2023); Zan et al. (2023). Despite this impressive progress, the performance of current open-source models still lags behind that of proprietary ones (e.g., GPT-3.5, GPT-4, etc.), primarily because stronger models often keep their training data proprietary Hui et al. (2024). As a result, the lack of publicly available code datasets remains a significant barrier to further development in this field.
@@ 3 WarriorCoder: Learning from Expert Battles
Figure 2: The diagram of learning from expert battles. In each round of the arena, the attacker challenges the defender in its area of expertise under the evaluation of judges, and then the winner’s response is added to the training data. In this manner, the target model gradually incorporates the strengths of all the code experts by fine-tuning on the data.
In this section, we describe how WarriorCoder learns from expert battles. Unlike previous approaches that expand existing datasets by prompting a limited pool of proprietary LLMs, we construct an arena where state-of-the-art code LLMs compete against each other. Each model leverages its learned knowledge to challenge others, while judges evaluate the outcomes. The target model then learns from the winner of these pairwise competitions, progressively integrating the strengths of all competitors.
3.1 Competitors Setting
The capabilities of competitors determine the final performance of WarriorCoder. Theoretically, the more diverse and high-quality training data are derived from a larger and stronger pool of competitors. For this study, we select five leading open-source code experts from the BigCodeBench Leaderboard Zhuo et al. (2024) - Athene-V2-Chat Su et al. (2025), DeepSeek-Coder-V2-Lite-Instruct DeepSeek-AI et al. (2024), Llama-3.3-70B-Instruct Dubey et al. (2024), Qwen2.5-72B-Instruct Hui et al. (2024), and QwQ-32B-Preview Team (2024). Notably, while WarriorCoder achieves state-of-the-art performance based solely on open-source code LLMs, it can also learn from powerful proprietary LLMs. In each round of the arena, only one pair of code experts is selected as competitors, while the remaining ones serve as judges.
@@ 3.2 Instruction Mining from Scratch
Considering a battle between LLM A and LLM B where A is the attacker and B is the defender. The first step of the arena is to use the strengths of A to challenge B, which makes it necessary to know what A has learned during its training process. However, almost all open-source LLMs keep their core data proprietary. Inspired by Magpie Xu et al. (2024b), we design a completion-based method to mine the capabilities which the code LLMs have already mastered (① Completion-based Instruction Mining). Here we take Qwen2.5 Hui et al. (2024) as an example. A conversation about writing Python code in the chat template of Qwen2.5 is:
@@
Such chat templates are pre-defined conversational structures to guide the interaction between the model and the user. Based on the LLMs’ strong completion abilities, we feed only the prefix of the chat template into them, prompting the LLM to generate the user instructions:
@@
In this way, we can collect various instructions $I$ the model has already learned under various generation settings (different values of temperature and top-p). Unlike traditional data synthesis, $I$ is not synthesized by the models but directly sampled from their distributions, which avoids pattern overfitting and significant shifts in the output distribution Chen et al. (2024b). However, these instructions may be repetitive, ambiguous, unclear, or too easy. To address these concerns, we deduplicate the data and adopt judges to assess their difficulty (② Deduplication and Difficulty Filtering). We divide the difficulty of instructions into four levels:
  • Excellent (9-10): For instructions that are very clear, specific, and well-articulated. These instructions are particularly challenging and excellently designed to assess the AI’s proficiency.
  • Good (6-8): For instructions that are clear and specific instructions. These are not overly difficult to answer and moderately assess the AI’s capabilities.
  • Average (3-5): For instructions that are fairly clear and specific instructions. These instructions are easy to answer.
  • Poor (1-2): For instructions that are ambiguous or unclear.
Only good and excellent instructions are considered during the following steps:
$\bar{I}=\{i|i\in I\wedge d(i)\geq 6\}$ (1)
where $d(i)$ is the difficulty of instruction $i$.
Then we compress the high-quality instructions $\bar{I}$ for the efficiency of post-training (③ Embedding-based Compression). To ensure the diversity and representativeness of instructions, we employ KCenterGreedy algorithm Sener and Savarese (2018) to select the final instructions $\bar{\bar{I}}$ based on the embedding model - all-roberta-large-v1 Liu et al. (2019).
@@ 3.3 Win-Loss Decision
The defender is required to respond to the attacker’s question, while the attacker A must also provide an answer to its own instruction (④ Code Generating). Once both answers are collected, the judges (the rest LLMs in arena) will evaluate the correctness and helpfulness of the pairwise responses and vote for their preferred one (more details can be found in Appendix C). Then we can calculate the local score for each response:
$x^{i}_{A>B}=\frac{t_{A}}{t_{A}+t_{B}}\quad x^{i}_{B>A}=\frac{t_{B}}{t_{A}+t_{B}}$ (2)
where $x^{i}_{A>B}$ and $x^{i}_{B>A}$ are the local scores for A’s and B’s responses to the instruction $i$. $x^{i}_{A>B}$ represents the percentage of votes that candidate A receives, while $x^{i}_{B>A}$ similarly represents the percentage of votes that candidate B receives. $t_{A}$ and $t_{B}$ are the number of votes which A and B win.
@@
where $X^{Elo}_{A>B}$ and $X^{Elo}_{B>A}$ indicate the expected probabilities of A defeating B and B defeating A, respectively. $R_{A}$ and $R_{B}$ are the Elo rating of A and B, which are updated dynamically and iteratively. Given the battle result of A and B on an instruction $i$, we update them by:
$R_{A}\leftarrow R_{A}+K\times(s^{i}_{A>B}-X^{Elo}_{A>B})$ (4)
$R_{B}\leftarrow R_{B}+K\times(s^{i}_{B>A}-X^{Elo}_{B>A})$
where $s^{i}_{A>B}$ and $s^{i}_{B>A}$ are the actual score of the battle result of player A and B (1 for a win, 0.5 for a draw, and 0 for a loss). The factor $K$ controls the sensitivity of rating changes.
Based on Equation 2 and Equation 3, we can obtain the final score of A’s response for instruction $i$:
$e^{i}_{A}=\sum_{B\in Com\setminus A}\alpha X^{Elo}_{A>B}+(1-\alpha)x^{i}_{A>B}$ (5)
where $Com$ is the set of all the competitors and $\setminus$’ is the subtraction operation. $\alpha$ is the coefficient to balance the local contingency and global consistency.
@@ 3.4 Final Training
Each item in the constructed dataset consists of an instruction, responses from various strong LLMs, and their corresponding scores, which supports multiple post-training methods. We select the response with the highest score as the gold output to obtain Instructions with Best Responses in Figure 2 and use SFT to train the target model. In this manner, WarriorCoder integrates the strengths of all the code experts, as their expertise is embedded in the instructions and responses within the training data.
@@ 4 Experiments
Models HumanEval HumanEval+ MBPP MBPP+ Rely on proprietary LLMs?
@@
Ours WarriorCoder (6.7B) 80.5 (+32.9) 75.6 (+36.0) 76.2 (+6.0) 64.8 (+8.2)
Table 1: The pass@1(%) results on the code generation benchmarks (Humaneval, Humaneval+, MBPP and MBPP+).
4.1 Experimental Details
Backbones
We use DeepSeekCoder-Base-6.7B Guo et al. (2024) to initialize WarriorCoder. As for the competitors of expert battles, we choose strong open-source LLMs including Athene-V2-Chat Su et al. (2025), DeepSeek-Coder-V2-Lite-Instruct DeepSeek-AI et al. (2024), Llama-3.3-70B-Instruct Dubey et al. (2024), Qwen2.5-72B-Instruct Hui et al. (2024), and QwQ-32B-Preview Team (2024).
Datasets
To evaluate the code generation capability of WarriorCoder, we conduct evaluations on HumanEval Chen et al. (2021), MBPP Austin et al. (2021), HumanEval+ Liu et al. (2023), and MBPP+ Liu et al. (2024). Besides, we also evaluate its code reasoning and libraries usage capabilities on CRUXEval Gu et al. (2024) and DS-1000 Lai et al. (2022). For a fair comparison, we use the same decoding strategies and generation configs as the previous work Wei et al. (2024); Luo et al. (2024b); Yu et al. (2024).
Experimental Settings
During the intruction minging, we adopt 9 different generation configs where temperature $t\in\{1.0,1.1,1.2\}$ and top-p $p\in\{0.99,0.995,1.0\}$. The final number of battle rounds is set to 70,000 and $K$ is set to 40. $\alpha$ is set to 0.7 because we need the Elo Rating only when judges’ opinions are divided on the evaluation. The detailed prompts can be found in Appendix B. As for the training stage, we conduct parallel training on 8 NVIDIA A800 80G GPUs. The global batch size is set to 512, and the number of total training steps is set to 448. We use a learning rate of $1\times 10^{-5}$ and a weight decay of $3\times 10^{-7}$. Additionally, a WarmupLR scheduler with a warmup ratio of 0.2 is used.
@@ 4.2 Main Results
Models CRUXEval-I CRUXEval-O
Pass@1 Pass@5 Pass@1 Pass@5
Proprietary Models GPT-4-Turbo 69.8 76.8 68.7 73.0
GPT-3.5-Turbo 49.0 63.2 49.4 59.3
Claude-3-Opus 64.2 - 65.8 -
Open-source Models StarCoder (6.7B) 29.7 47.3 32.2 44.9
StarCoder (15B) 31.3 49.2 34.2 47.1
DeepSeekCoder-Instruct (6.7B) 37.4 53.3 41.2 52.8
CodeLlama-Python (6.7B) 37.3 57.0 35.9 48.8
CodeLlama-Python (13B) 39.7 56.9 39.8 52.5
CodeLlama-Python (34B) 43.9 59.5 41.4 52.9
Mistral (6.7B) 35.0 52.3 34.3 48.6
WizardCoder (13B) 36.5 51.6 41.3 52.4
WizardCoder (34B) 42.7 57.5 43.4 53.8
Magicoder(6.7B) 41.7 62.4 44.4 57.5
Ours WarriorCoder (6.7B) 42.9 66.5 45.4 66.3
Table 2: The pass@1(%) and pass@5(%) results on the code reasoning benchmark (CRUXEval).
Models Matplotlib NumPy Pandas PyTorch SciPy Sklearn TensorFlow Overall
(155) (220) (291) (68) (106) (115) (45) (1000)
INCODER (6.7B) 28.3 4.4 3.1 4.4 2.8 2.8 3.8 7.4
CodeGen-Mono (16B) 31.7 10.9 3.4 7.0 9.0 10.8 15.2 11.7
Code-Cushman-001 40.7 21.8 7.9 12.4 11.3 18.0 12.2 18.1
StarCoder (15B) 51.7 29.7 11.4 21.4 20.2 29.5 24.5 26.0
WizardCoder-SC (15B) 55.2 33.6 16.7 26.2 24.2 24.9 26.7 29.2
CodeLlama-Python (6.7B) 55.3 34.5 16.4 19.9 22.3 17.6 28.5 28.0
WizardCoder-CL (6.7B) 53.5 34.4 15.2 25.7 21.0 24.5 28.9 28.4
Magicoder-CL (6.7B) 54.6 34.8 19.0 24.7 25.0 22.6 28.9 29.9
MagicoderS-CL (6.7B) 55.9 40.6 28.4 40.4 28.8 35.8 37.6 37.5
WarriorCoder (6.7B) 55.5 41.8 26.1 41.2 33.0 39.1 42.2 38.1
Table 3: The pass@1(%) results on the benchmark for using Python libraries in data science (DS-1000).
Task Percentage(%) Definition
Code Generation 51.4 Generating source code based on certain specifications or requirements.
Code Debugging 12.2 Identifying, diagnosing, and fixing errors or bugs in a code snippet.
Code Optimization 3.8 Improving a program’s performance, efficiency, or resource usage without changing its functionality.
Code Reasoning 2.9 Predicting the output based on the given input or predicting the input from the known output.
Code Analysis 6.6 Analyzing, understanding, and explaining how a piece of code works.
Theoretical Explanation 22.2 Answering the questions about principles, theories, and properties of programming language.
Code Transpile 0.9 Converting source code from one programming language into another programming language.
Table 4: The proportion of different tasks in the training data.
@@
The results on the code generation benchmarks are summarized in Table 1. WarriorCoder achieves SOTA performance, with a pass@1 accuracy of 80.5% (75.6%) in HumanEval (HumanEval +) and 76.2% (64.8%) in MBPP (MBPP +), surpassing all other fine-tuned models. Particularly, it shows a significant boost over on HumanEval and HumanEval+ (with gains of 32.9 and 36.0, respectively). WarriorCoder also outperforms MagicoderS-DS, Magicoder-DS and WaveCoder-DS, which share the same backbone architecture and similar amounts of training data. This substantial performance gap highlights the effectiveness of our approach in generating higher-quality training data, providing a clear advantage over models that rely on similar foundational setups.
Moreover, WarriorCoder also achieves excellent performances on the code reasoning benchmark and libraries usage benchmark. As shown in Table 2, WarriorCoder outperforms a range of open-source models, including those with sizes up to 34B, in pass@1 accuracy and achieves better pass@5 accuracy compared to GPT-3.5-Turbo (66.5% vs 63.2% on CRUXEval-I and 66.3% vs 59.3% on CRUXEval-O). Table 3 shows that WarriorCoder outperforms all baselines on most of the libraries, especially on SciPy, Sklearn and Tensorflow (33.0% , 39.1% , and 42.2%, respectively). These results highlight WarriorCoder as a powerful paradigm - a data flywheel that absorbs expertise from multiple code domains. Our approach significantly enhances the target model’s ability to generalize across various tasks, demonstrating its superiority in leveraging diverse data sources to drive performance improvements.
@@
Addiontionally, previous data flywheels typically rely on augmentations and annotations generated using proprietary LLMs and specially designed prompts. In contrast, our approach does not require pre-existing datasets, diverse handwritten prompts or proprietary LLMs. Despite this, we get competitive results that rival those of advanced proprietary code experts. This highlights the effectiveness of our data flywheel, demonstrating the feasibility of collecting high-quality data at a low cost.